# Detection of very high-energy gamma-ray emission from the radio galaxy M87 with LHAASO

### v1

@@ 1 Introduction
The radio galaxy M87 is one of the nearest active galactic nuclei (AGN) located in the Virgo Cluster at a distance of approximately $16.8\pm 0.8$ Mpc (Blakeslee et al., 2009). It hosts a super massive black hole (SMBH), known as M87, with a mass of $(\rm 6.5\pm 0.2_{stat}\pm 0.7_{sys})\times 10^{9}~{}M_{\odot}$ (Event Horizon Telescope Collaboration et al., 2019). M87 has been a subject of extensive studies across the entire electromagnetic spectrum (see EHT MWL Science Working Group et al. (2021) for a review of recent observations).
@@
The first evidence for very-high-energy (VHE) gamma-ray emission from M87 was reported by the High Energy Gamma Ray Astronomy during 1998-1999 (Aharonian et al., 2003). It was later confirmed by the next-generation imaging air Cherenkov telescopes (IACTs), including H.E.S.S. (Aharonian et al., 2006), VERITAS (Acciari et al., 2009, 2010), MAGIC (Albert et al., 2008), as well as by the High Altitude Water Cherenkov Observatory (HAWC) at a marginal statistical level (Albert et al., 2021). M87 has also been detected by Fermi-LAT in the high-energy (HE, $0.1-30{~{}\rm GeV}$) band (Abdo et al., 2009).
@@
The exact location of the VHE gamma-ray emitting region in M87 remains unknown. The angular resolution of ground-based VHE instruments is of the order of 0.1 degree (corresponding to 30 kpc projected size) and, therefore, does not allow to locate the emitting region. The suggested sites of TeV gamma-ray production range from the immediate environment (e.g., the magnetosphere) of the supermassive black hole(Aharonian et al., 2006; Neronov & Aharonian, 2007; Rieger & Aharonian, 2012) to inner (subparsec) parts of the jet (Reimer et al., 2004; Georganopoulos et al., 2005), a compact hot spot (the so-called HST-1 knot) at a distance of 100 pc along the jet (Stawarz et al., 2006), and large-scale structures in the kiloparsec jet (Stawarz et al., 2005). To further investigate the location of the VHE gamma-ray emission site, variability studies and the search for correlations with other wavelengths have usually been used (e.g., Acciari et al., 2009).
@@
Four major VHE flares from M87 have been reported before the current observations, in 2005 (Aharonian et al., 2006), 2008 (Albert et al., 2008; Acciari et al., 2009), 2010 (Abramowski et al., 2012) and 2018 (H. E. S. S. Collaboration et al., 2024; The Event Horizon Telescope-Multi-wavelength science working group et al., 2024). The variability time scales of the these VHE flares is about one day. This points, through the causality argument, to a compact emission region with a size smaller than $3\times 10^{15}\delta\,{\rm cm}$, where $\delta\sim{\rm a\,\,few}$ is the Doppler factor of the emitting region. The coincidence of the first VHE flare with a giant X-ray flare from HST-1 (Harris et al., 2006) initially prompted speculations that the observed VHE emission may have originated in HST-1 (e.g., Stawarz et al. (2006)). During the second flaring episode, however, HST-1 was in a low flux state, while radio measurements showed a flux increase in the core region within a few hundred Schwarzschild radii of the SMBH, suggesting the direct vicinity of the SMBH as the site of the VHE gamma-ray emission (Acciari et al., 2009). This conclusion was further supported by the detection of an enhanced X-ray flux from the core by Chandra. Therefore, the potential site of such rapid flares is related to either the core of M87 (Aharonian et al., 2006) or to very compact regions in the jet, for instance, the HST-1 knot (Harris et al., 2006). Note that, however, the location of HST-1, at a distance of 100 pc from the central BH, would require an extremely tight collimation of the jet to account for the observed fast (days-scale) TeV variability (Aharonian et al., 2006). Moreover, the fact that the Chandra X-ray flux from HST-1 did not change much during subsequent VHE flaring episodes (after the one in 2005) has been taken to disfavor HST-1 as a site of the rapid TeV flaring activity (Rieger & Aharonian, 2012).
@@
2 LHAASO observations
LHAASO is a multifaceted and comprehensive EAS detector array. The facility consists of three sub-arrays, including the Water Cherenkov Detector Array (WCDA, sensitive to VHE gamma-rays at energy from a few hundreds of GeV to tens of TeV), the 1.3 km2 ground array (KM2A, sensitive to VHE gamma-rays at energy above 20 TeV) and the Wide Field-of-view Air Cherenkov/Fluorescence Telescope Array (WFCTA), designed for cosmic ray studies from 10 TeV to 1 EeV. The detailed information about the performance of these sub-arrays and data reconstruction algorithms can be seen in Cao et al. (2019, 2024). The details of data analysis are summarized in Appendix A. The first source catalog of Large High Altitude Air Shower Observatory reported the detection of a very-high-energy gamma ray source, 1LHAASO J1219+2915 (Cao et al., 2024), which is identified as the counterpart of a low-luminosity AGN (NGC 4278 Cao et al., 2024b).
The data utilized for M87 in this work is collected by full array of LHAASO-WCDA from 8 March, 2021 to 16 March 2024 and full array of LHAASO-KM2A from 20 July, 2021 to 31 January, 2024. The effective live time is 1026 days and 884 days, respectively. The number of fired hit ($N_{\rm{hit}}$) is selected as the estimator of primary energy for WCDA (Aharonian et al., 2021a). The events are binned into seven segments with $N_{\rm{hit}}$ value of [60, 100), [100, 200), [200, 300), [300, 500), [500, 700), [700, 1000) and [1000, 2000]. The KM2A events are divided into bins with reconstructed energy width of 0.2 in log scale (Aharonian et al., 2021b). The “direct integral method” is adopted here to estimated cosmic ray background (Fleysher et al., 2004). A binned maximum likelihood analysis based on the forward folding method ( i.e., convolving the assumed spectrum model with the detector response) is performed to estimate best-fit values of position and spectrum parameters. The test statistic (TS) is used to calculate significance of the target source, $\rm{TS}=-2(\ln\mathcal{L}_{\rm{b}}-\ln\mathcal{L}_{\rm{s+b}})$, where $\mathcal{L}_{\rm{b}}$ and $\mathcal{L}_{\rm{s+b}}$ are the likelihood values for the background without (null hypothesis) and with the target signal model, respectively.
@@ 2 LHAASO observations
2.1 The full-time average spectrum
@@
The cutoff shape at the high energy in the observed SED could be partly or fully caused by the extra-galactic background light (EBL) absorption. The intrinsic spectrum can be obtained by correcting for the absorption due to EBL. The simplest assumption for the intrinsic spectrum is that it is described by a power-law function $dN/dE=N_{0}(E/{\rm 3TeV})^{-\gamma}$. To incorporate the EBL absorption into our spectral models, we select three models that represent the uncertain range of the EBL intensity, the kneiske model (Kneiske & Dole, 2010) for the lower boundary of the EBL flux, the franceschini model for the medium intensity (Franceschini et al., 2008), and dominguez-upper model reflecting upper bounds of the EBL flux (Domínguez et al., 2011). As demonstrated in Table 1, these PLxEBL models all yield good fit to the observed spectral data with slightly improved $\chi^{2}/{\rm ndf}$ compared with the PLEC and LP fit. A higher EBL flux results in a harder intrinsic spectrum, although the statistical error in the photon index is still large. For subsequent analysis, we adopt the medium-intensity EBL model (the franceschini model) as the baseline EBL model. The photon attenuation derived from the EBL model as a function of energy at the distance of M87 (z = 0.0042) is applied to correct the detection efficiency. With EBL-corrected detection efficiency, the photon index of the intrinsic spectrum is 2.37$\pm$0.14 for the full-time period. The best-fit parameters for the power-law intrinsic spectrum are summarized in Table 2. As shown in the left panel of Figure 3, the intrinsic spectrum for the full time data extends to tens of TeV without a visible softening. Interestingly, a possible spectral hardening is seen at 20 TeV.
@@ 2.2 The light curve
For the two-day binning light curve, the likelihood variability test (Abeysekara et al., 2017) of a constant flux (set to the average flux of full time) gives a p-value of $4\times 10^{-3}$ ($\sim 2.7\sigma$), indicating the possible existence of variability. In order to quantify the fluctuations of temporal profiles, we employ the Bayesian block method (Scargle et al., 2013) on the two-day binning light curve with a prior false positive probability of 5%. Only one clear block from MJD 59607 to MJD 59615 is found, which is shown in the upper-right panel of Figure 4 (the blue line). We identify this period as a flare, whose details will be discussed in section 2.2.1. Four time window widths of 1 day, 2 days, 4 days, and 8 days were used to search the 1026-day LHAASO-WCDA data, resulting in 2071 searches. A conservative way is adopted to calculate the post-trial significance of the flare by setting the trial number equal to 2071, which yields a post-trial significance of $\sim 4.4\,\sigma$.
2.2.1 The 2022 VHE flare
The best-fit position of the emission detected during the flare period is (R.A. =$187.75^{\circ}\pm 0.06^{\circ}$, Decl. = $12.39^{\circ}\pm 0.06^{\circ}$), which is 0.07 away from radio position of core of M87, as shown in the middle panel of Figure 1. The TS value and photon index of the intrinsic spectrum for this flare are 40.1 and $2.57\pm 0.23$ (see Table 2), respectively. The spectrum of the flare is also shown in Figure 3 (the middle panel).
The VHE flare detected by LHAASO has an average flux of $(3.78\pm 0.71)\times 10^{-12}~{}\rm photons~{}cm^{-2}~{}s^{-1}$, corresponding to an energy flux of $1.38\times 10^{-11}{\rm~{}erg~{}cm^{-2}~{}s^{-1}}$ in 1-20 TeV assuming a photon index $\gamma=2.57\pm 0.23$. With a distance of $d=16.8\,\mathrm{Mpc}$, the corresponding isotropic VHE luminosity is $\sim 4.7\times 10^{41}~{}{\rm erg~{}s^{-1}}$. This is a non-negligible amount compared to the bolometric luminosity ($\sim 10^{42}\,\mathrm{erg\,s^{-1}}$) of the M87 nucleus (Kharb & Shastri, 2004) and the total kinetic luminosity ($\sim 10^{44}{\rm erg~{}s^{-1}}$) of the jet (Owen et al., 2000).
@@ 2.2.1 The 2022 VHE flare
$F=F_{0}\times e^{-|t-t_{0}|/\Delta\tau}{\rm with}\left\{\begin{aligned} \Delta\tau&=\Delta\tau^{\rm rise}&\textrm{for}\enspace t<t_{0}\\ \Delta\tau&=\Delta\tau^{\rm decay}&\textrm{for}\enspace t\geq t_{0}.\end{aligned}\right.$ (3)
The peak time $t_{0}$ is set to be free, and the resulting flux doubling time $\tau_{\rm d}={\rm ln(2)}\times\Delta\tau$ of the rising part is $\tau_{\rm d}^{\rm rise}=1.05\pm 0.49$ days and that of the decaying part is $\tau_{\rm d}^{\rm decay}=2.17\pm 0.58$ days. The variability time scale is similar to that of the 2010 flare (Abramowski et al., 2012), but the 2022 flare reported in this work rises faster and decays slower. The fitting result is shown in Figure 5. This obtained variability time scale $\tau_{d}=\tau_{d}^{\rm rise}=1.05\pm 0.49$ days corresponds to the size of a spherical emission region of $R\leq c\tau_{d}\delta\simeq 2.7\times 10^{15}\delta~{}\rm cm$, where $\delta$ is the Doppler factor to describe the relativistic beaming.
2.2.2 The low-state emission
We consider the non-flare period (i.e., excluding the flare period from MJD 59607 to MJD 59615) as the low-state. The significance map of this low-state emission is shown in the right panel of Figure 1 and the best-fit position, marked by the cyan open cross, is consistent with the radio position of the core of M87. The TS value of the low-state emission is 76.0. The photon index of the low-state emission, 2.31$\pm$0.17, is similar to the one obtained with fitting the entire data set. This photon index is consistent with that of the low state emission of M87 measured by H.E.S.S. (H. E. S. S. Collaboration et al., 2023). The spectral data for the low-state emission is shown in Figure 3 (the right panel).
3 Fermi-LAT observations
@@
The Fermi-LAT data (Atwood et al., 2009) of M87 are extracted from Fermi Science Support Center [^1] https://fermi.gsfc.nasa.gov/ssc/data/access/. The data cover about 15.5 years from August 4, 2008 to March 16, 2024, in an energy range between 100 MeV and 100 GeV. The details of analysis are summarized in Appendix B. For the 15.5-year observation of M87, the source is detected with TS=2434.31 (corresponding to a significance of $\sim 49\sigma$). The average flux is $F=(1.79\pm 0.10)\times 10^{-8}~{}\rm photons~{}cm^{-2}~{}s^{-1}$, with a photon index of $\Gamma=2.05\pm 0.03$. The average flux is shown by the gray lines in the bottom panels of Figure 4.
@@
4 Discussions
4.1 VHE flare emission
@@
For the variability time of $\tau=1.05$ days of the current VHE flare, the size of the emission region should be smaller than $2.7\times 10^{15}\delta\,{\rm cm}$, where $\delta$ is the Doppler factor of the emitting region. The jet of M87 is observed at a large angle $\theta\sim 20$ degrees (Biretta et al., 1999), making M87 a “misaligned” BL Lacertae object that most jet emission is characterized by rather weak Doppler factors $\delta=1/[\Gamma_{j}(1-\beta_{j}\cos\theta)]\lesssim 3$, where $\Gamma_{j}\sim 10$ is the jet Lorentz factor. Since the mass of the SMBH in M87 is well established, $M=6\times 10^{9}M_{\odot}$, the Schwarzschild radius is estimated quite accurately: $R_{s}=2GM/c^{2}\simeq 2\times 10^{15}{\rm cm}$. Therefore, the size of the emission region of the VHE flare is only a few Schwarzschild radii of the SMBH.
@@
The small size of the TeV emission region, together with the correlated variability between the X-ray emission (and radio emission) of the core and the TeV emission in some previous VHE flares, suggest that the relevant non-thermal particle acceleration and emission processes could take place in the core of M87. One scenario involves inner sub-pc jets that produce VHE emission like blazars (Georganopoulos et al., 2005; Tavecchio & Ghisellini, 2008). Georganopoulos et al. (2005) considered the case of a relativistic jet decelerating substantially on sub-parsec distances from the core, and showed that the velocity difference between a faster and a slower portion of the outflow could lead to the enhanced inverse-Compton emission of the former one in the TeV range.
@@
Magnetospheric particle acceleration and emission models have also been invoked to explain the observed VHE emission in M87 (Neronov & Aharonian, 2007; Rieger & Aharonian, 2008). Such a mechanism of gamma-ray emission from the vicinity of a BH is a close analog of the mechanism of pulsed gamma-ray emission from the vicinity of neutron stars in pulsars. Usually, efficient particle acceleration in these scenarios is related either to gap-type (Neronov & Aharonian, 2007; Levinson & Rieger, 2011) or centrifugal-type processes occurring in the magnetosphere around a rotating black hole (Rieger & Aharonian, 2008). Because of its low infrared luminosity, the nucleus of M87 can be effectively transparent for gamma-rays up to an energy of $\sim 10$ TeV.
@@
Other types of models have been also put forward to explain the observed rapid variability of the VHE $\gamma$-ray emission from M87. Recent high-resolution general-relativistic magnetohydrodynamic simulations show the occurrence of episodic magnetic reconnection events that can power flares near the black hole event horizon (Ripperda et al., 2020; Hakobyan et al., 2023). It was also argued that interactions of a relativistic magnetized outflow around the jet formation zone with a star partially tidally disrupted by the interaction with the SMBH can lead to an efficient production of VHE photons via hadronic processes (Barkov et al., 2010; Aharonian et al., 2017).
@@
It is worth to mention that that, unlike IACTs, LHAASO data provide continuous and uniform TeV measurements of M87. While IACTs observations have a better sensitivity for short-time exposure, LHAASO observations provide a comprehensive study of the TeV light curve on the months and years time scales. The measurements presented in this work can be used to estimate the duty cycle of VHE flares in M87. The duty cycle of astrophysical sources is defined as the fraction of time during which the sources are in a flaring state. This can be used to estimate the active time of the central engine and the total energy release during the flaring state, in comparison with that in the low-state. Previous IACT observations of M87 found that the respective duty cycle of VHE flares are $\sim 14\%$, $7\%$ and $4\%$ for a threshold flux of $0.5\times 10^{-11}{\rm cm^{-2}s^{-1}}$, $0.8\times 10^{-11}{\rm cm^{-2}s^{-1}}$ and $10^{-11}{\rm cm^{-2}s^{-1}}$(Abramowski et al., 2012). However, due to the observing strategy for M87, i.e., observations have been intensified after the detection of a high state, the data set could be biased. The derived duty cycle is, therefore, likely overestimated and should be considered an upper limit(Abramowski et al., 2012). During the three-year continuous monitoring of M87, LHAASO detected one strong flare over a period of 8 days with an energy flux reaching $1.38\times 10^{-11}{\rm~{}erg~{}cm^{-2}~{}s^{-1}}$ in 1–20 TeV, revealing a duty cycle of $\sim{\rm 8d/3yr}\sim 1\%$ for such bright flares. As the flux during the flaring state is one order of magnitude of higher, the total energy release during the flaring state is about 10% of the energy release in the low state. This may be useful to constrain the underlying physical process for VHE flares.
4.2 Long-term low-state VHE emission

### v3

@@ I Introduction
The radio galaxy M87 is one of the nearest active galactic nuclei (AGN) located in the Virgo Cluster at a distance of approximately $16.8\pm 0.8$ Mpc (2009ApJ...694..556B). It hosts a super massive black hole (SMBH), known as M87, with a mass of $(\rm 6.5\pm 0.2_{stat}\pm 0.7_{sys})\times 10^{9}~M_{\odot}$ (2019ApJ...875L...6E). M87 has been a subject of extensive studies across the entire electromagnetic spectrum (see 2021ApJ...911L..11E for a review of recent observations).
@@
The first evidence for very-high-energy (VHE) gamma-ray emission from M87 was reported by the High Energy Gamma Ray Astronomy during 1998-1999 (2003A&A...403L...1A). It was later confirmed by the next-generation imaging air Cherenkov telescopes (IACTs), including H.E.S.S. (2006Sci...314.1424A), VERITAS (2009Sci...325.. [^4] A; 2010ApJ...716..819A), MAGIC (2008ApJ...685L..23A), as well as by the High Altitude Water Cherenkov Observatory (HAWC) at a marginal statistical level (2021ApJ...907...67A). M87 has also been detected by Fermi-LAT in the high-energy (HE, $0.1-30{~\rm GeV}$) band (2009ApJ...707...55A).
@@
The exact location of the VHE gamma-ray emitting region in M87 remains unknown. The angular resolution of ground-based VHE instruments is of the order of 0.1 degree (corresponding to 30 kpc projected size) and, therefore, does not allow to locate the emitting region. The suggested sites of TeV gamma-ray production range from the immediate environment (e.g., the magnetosphere) of the supermassive black hole(2006Sci...314.1424A; 2007ApJ...671...85N; 2012MPLA...2730030R) to inner (subparsec) parts of the jet (2004A&A...419...89R; 2005ApJ...634L..33G), a compact hot spot (the so-called HST-1 knot) at a distance of 100 pc along the jet (2006MNRAS.370..981S), and large-scale structures in the kiloparsec jet (2005ApJ...626..120S). To further investigate the location of the VHE gamma-ray emission site, variability studies and the search for correlations with other wavelengths have usually been used (e.g., 2009Sci...325.. [^4] A).
@@
Four major VHE flares from M87 have been reported before the current observations, in 2005 (2006Sci...314.1424A), 2008 (2008ApJ...685L..23A; 2009Sci...325.. [^4] A), 2010 (2012ApJ...746..151A) and 2018 (2024A&A...685A..96H; 2024arXiv240417623T). The variability time scales of the these VHE flares is about one day. This points, through the causality argument, to a compact emission region with a size smaller than $3\times 10^{15}\delta\,{\rm cm}$, where $\delta\sim{\rm a\,\,few}$ is the Doppler factor of the emitting region. The coincidence of the first VHE flare with a giant X-ray flare from HST-1 (2006ApJ...640..211H) initially prompted speculations that the observed VHE emission may have originated in HST-1 (e.g., 2006MNRAS.370..981S). During the second flaring episode, however, HST-1 was in a low flux state, while radio measurements showed a flux increase in the core region within a few hundred Schwarzschild radii of the SMBH, suggesting the direct vicinity of the SMBH as the site of the VHE gamma-ray emission (2009Sci...325.. [^4] A). This conclusion was further supported by the detection of an enhanced X-ray flux from the core by Chandra. Therefore, the potential site of such rapid flares is related to either the core of M87 (2006Sci...314.1424A) or to very compact regions in the jet, for instance, the HST-1 knot (2006ApJ...640..211H). Note that, however, the location of HST-1, at a distance of 100 pc from the central BH, would require an extremely tight collimation of the jet to account for the observed fast (days-scale) TeV variability (2006Sci...314.1424A). Moreover, the fact that the Chandra X-ray flux from HST-1 did not change much during subsequent VHE flaring episodes (after the one in 2005) has been taken to disfavor HST-1 as a site of the rapid TeV flaring activity (2012MPLA...2730030R).
@@
II LHAASO observations
LHAASO is a multifaceted and comprehensive EAS detector array. The facility consists of three sub-arrays, including the Water Cherenkov Detector Array (WCDA, sensitive to VHE gamma-rays at energy from a few hundreds of GeV to tens of TeV), the 1.3 km2 ground array (KM2A, sensitive to VHE gamma-rays at energy above 20 TeV) and the Wide Field-of-view Air Cherenkov/Fluorescence Telescope Array (WFCTA), designed for cosmic ray studies from 10 TeV to 1 EeV. The detailed information about the performance of these sub-arrays and data reconstruction algorithms can be seen in 2019arXiv190502773C; 2024ApJS..271...25C. The details of data analysis are summarized in Appendix A. The first source catalog of Large High Altitude Air Shower Observatory reported the detection of a very-high-energy gamma ray source, 1LHAASO J1219+2915 (2024ApJS..271...25C), which is identified as the counterpart of a low-luminosity AGN (NGC 4278 Cao_2024b).
The data utilized for M87 in this work is collected by full array of LHAASO-WCDA from 8 March, 2021 to 16 March 2024 and full array of LHAASO-KM2A from 20 July, 2021 to 31 January, 2024. The effective live time is 1026 days and 884 days, respectively. The number of fired hit ($N_{\rm{hit}}$) is selected as the estimator of primary energy for WCDA (Aharonian_2021a). The events are binned into seven segments with $N_{\rm{hit}}$ value of [60, 100), [100, 200), [200, 300), [300, 500), [500, 700), [700, 1000) and [1000, 2000]. The KM2A events are divided into bins with reconstructed energy width of 0.2 in log scale (Aharonian_2021b). The “direct integral method” is adopted here to estimated cosmic ray background (Fleysher_2004). A binned maximum likelihood analysis based on the forward folding method ( i.e., convolving the assumed spectrum model with the detector response) is performed to estimate best-fit values of position and spectrum parameters. The test statistic (TS) is used to calculate significance of the target source, $\rm{TS}=-2(\ln\mathcal{L}_{\rm{b}}-\ln\mathcal{L}_{\rm{s+b}})$, where $\mathcal{L}_{\rm{b}}$ and $\mathcal{L}_{\rm{s+b}}$ are the likelihood values for the background without (null hypothesis) and with the target signal model, respectively.
@@ II LHAASO observations
II.1 The full-time average spectrum
@@
[…]
@@ II.2 The light curve
For the two-day binning light curve, the likelihood variability test (Abeysekara_2017) of a constant flux (set to the average flux of full time) gives a p-value of $4\times 10^{-3}$ ($\sim 2.7\sigma$), indicating the possible existence of variability. In order to quantify the fluctuations of temporal profiles, we employ the Bayesian block method (2013ApJ...764..167S) on the two-day binning light curve with a prior false positive probability of 5%. Only one clear block from MJD 59607 to MJD 59615 is found, which is shown in the upper-right panel of Figure 4 (the blue line). We identify this period as a flare, whose details will be discussed in section II.2.1. Four time window widths of 1 day, 2 days, 4 days, and 8 days were used to search the 1026-day LHAASO-WCDA data, resulting in 2071 searches. A conservative way is adopted to calculate the post-trial significance of the flare by setting the trial number equal to 2071, which yields a post-trial significance of $\sim 4.4\,\sigma$.
II.2.1 The 2022 VHE flare
The best-fit position of the emission detected during the flare period is (R.A. $187.75^{\circ}\pm 0.06^{\circ}$, Decl. = $12.39^{\circ}\pm 0.06^{\circ}$), which is 0.07 away from radio position of core of M87, as shown in the middle panel of Figure 1. The TS value and photon index of the intrinsic spectrum for this flare are 40.1 and $2.57\pm 0.23$ (see Table 2), respectively. The spectrum of the flare is also shown in Figure 3 (the middle panel).
The VHE flare detected by LHAASO has an average flux of $(3.78\pm 0.71)\times 10^{-12}~\rm photons~cm^{-2}~s^{-1}$, corresponding to an energy flux of $1.38\times 10^{-11}{\rm~erg~cm^{-2}~s^{-1}}$ in 1-20 TeV assuming a photon index $\gamma=2.57\pm 0.23$. With a distance of $d=16.8\,\mathrm{Mpc}$, the corresponding isotropic VHE luminosity is $\sim 4.7\times 10^{41}~{\rm erg~s^{-1}}$. This is a non-negligible amount compared to the bolometric luminosity ($\sim 10^{42}\,\mathrm{erg\,s^{-1}}$) of the M87 nucleus (2004A&A...425..825K) and the total kinetic luminosity ($\sim 10^{44}{\rm erg~s^{-1}}$) of the jet (2000ApJ...543..611O).
@@ II.2.1 The 2022 VHE flare
$F=F_{0}\times e^{-|t-t_{0}|/\Delta\tau}{\rm with}\left\{\begin{aligned} \Delta\tau&=\Delta\tau^{\rm rise}&\textrm{for}\kern 4.62497ptt<t_{0}\\ \Delta\tau&=\Delta\tau^{\rm decay}&\textrm{for}\kern 4.62497ptt\geq t_{0}.\end{aligned}\right$. (3)
The peak time $t_{0}$ is set to be free, and the resulting flux doubling time $\tau_{\rm d}={\rm ln(2)}\times\Delta\tau$ of the rising part is $\tau_{\rm d}^{\rm rise}=1.05\pm 0.49$ days and that of the decaying part is $\tau_{\rm d}^{\rm decay}=2.17\pm 0.58$ days. The variability time scale is similar to that of the 2010 flare (2012ApJ...746..151A), but the 2022 flare reported in this work rises faster and decays slower. The fitting result is shown in Figure 5. This obtained variability time scale $\tau_{d}=\tau_{d}^{\rm rise}=1.05\pm 0.49$ days corresponds to the size of a spherical emission region of $R\leq c\tau_{d}\delta\simeq 2.7\times 10^{15}\delta~\rm cm$, where $\delta$ is the Doppler factor to describe the relativistic beaming.
II.2.2 The low-state emission
[…]
III Fermi-LAT observations
@@
The Fermi-LAT data (2009ApJ...697.1071A) of M87 are extracted from Fermi Science Support Center [^1] https://fermi.gsfc.nasa.gov/ssc/data/access/. The data cover about 15.5 years from August 4, 2008 to March 16, 2024, in an energy range between 100 MeV and 100 GeV. The details of analysis are summarized in Appendix B. For the 15.5-year observation of M87, the source is detected with TS=2434.31 (corresponding to a significance of $\sim 49\sigma$). The average flux is $F=(1.79\pm 0.10)\times 10^{-8}~\rm photons~cm^{-2}~s^{-1}$, with a photon index of $\Gamma=2.05\pm 0.03$. The average flux is shown by the gray lines in the bottom panels of Figure 4.
@@
IV Discussions
IV.1 VHE flare emission
@@
For the variability time of $\tau=1.05$ days of the current VHE flare, the size of the emission region should be smaller than $2.7\times 10^{15}\delta\,{\rm cm}$, where $\delta$ is the Doppler factor of the emitting region. The jet of M87 is observed at a large angle $\theta\sim 20$ degrees (1999ApJ...520..621B), making M87 a “misaligned” BL Lacertae object that most jet emission is characterized by rather weak Doppler factors $\delta=1/[\Gamma_{j}(1-\beta_{j}\cos\theta)]\lesssim 3$, where $\Gamma_{j}\sim 10$ is the jet Lorentz factor. Since the mass of the SMBH in M87 is well established, $M=6\times 10^{9}M_{\odot}$, the Schwarzschild radius is estimated quite accurately: $R_{s}=2GM/c^{2}\simeq 2\times 10^{15}{\rm cm}$. Therefore, the size of the emission region of the VHE flare is only a few Schwarzschild radii of the SMBH.
@@
The small size of the TeV emission region, together with the correlated variability between the X-ray emission (and radio emission) of the core and the TeV emission in some previous VHE flares, suggest that the relevant non-thermal particle acceleration and emission processes could take place in the core of M87. One scenario involves inner sub-pc jets that produce VHE emission like blazars (2005ApJ...634L..33G; 2008MNRAS.385L..98T). 2005ApJ...634L..33G considered the case of a relativistic jet decelerating substantially on sub-parsec distances from the core, and showed that the velocity difference between a faster and a slower portion of the outflow could lead to the enhanced inverse-Compton emission of the former one in the TeV range.
@@
Magnetospheric particle acceleration and emission models have also been invoked to explain the observed VHE emission in M87 (2007ApJ...671...85N; 2008A&A...479L...5R). Such a mechanism of gamma-ray emission from the vicinity of a BH is a close analog of the mechanism of pulsed gamma-ray emission from the vicinity of neutron stars in pulsars. Usually, efficient particle acceleration in these scenarios is related either to gap-type (2007ApJ...671...85N; 2011ApJ...730..123L) or centrifugal-type processes occurring in the magnetosphere around a rotating black hole (2008A&A...479L...5R). Because of its low infrared luminosity, the nucleus of M87 can be effectively transparent for gamma-rays up to an energy of $\sim 10$ TeV.
@@
Other types of models have been also put forward to explain the observed rapid variability of the VHE $\gamma$-ray emission from M87. Recent high-resolution general-relativistic magnetohydrodynamic simulations show the occurrence of episodic magnetic reconnection events that can power flares near the black hole event horizon (2020ApJ...900..100R; 2023ApJ...943L..29H). It was also argued that interactions of a relativistic magnetized outflow around the jet formation zone with a star partially tidally disrupted by the interaction with the SMBH can lead to an efficient production of VHE photons via hadronic processes (2010ApJ...724.1517B; 2017ApJ...841...61A).
@@
It is worth to mention that that, unlike IACTs, LHAASO data provide continuous and uniform TeV measurements of M87. While IACTs observations have a better sensitivity for short-time exposure, LHAASO observations provide a comprehensive study of the TeV light curve on the months and years time scales. The measurements presented in this work can be used to estimate the duty cycle of VHE flares in M87. The duty cycle of astrophysical sources is defined as the fraction of time during which the sources are in a flaring state. This can be used to estimate the active time of the central engine and the total energy release during the flaring state, in comparison with that in the low-state. Previous IACT observations of M87 found that the respective duty cycle of VHE flares are $\sim 14\%$, $7\%$ and $4\%$ for a threshold flux of $0.5\times 10^{-11}{\rm cm^{-2}s^{-1}}$, $0.8\times 10^{-11}{\rm cm^{-2}s^{-1}}$ and $10^{-11}{\rm cm^{-2}s^{-1}}$(2012ApJ...746..151A). However, due to the observing strategy for M87, i.e., observations have been intensified after the detection of a high state, the data set could be biased. The derived duty cycle is, therefore, likely overestimated and should be considered an upper limit(2012ApJ...746..151A). During the three-year continuous monitoring of M87, LHAASO detected one strong flare over a period of 8 days with an energy flux reaching $1.38\times 10^{-11}{\rm~erg~cm^{-2}~s^{-1}}$ in 1–20 TeV, revealing a duty cycle of $\sim{\rm 8d/3yr}\sim 1\%$ for such bright flares. As the flux during the flaring state is one order of magnitude of higher, the total energy release during the flaring state is about 10% of the energy release in the low state. This may be useful to constrain the underlying physical process for VHE flares.
IV.2 Long-term low-state VHE emission